OpenStudy (anonymous):

george has 9 more dollars than Frank, who has 3 more than Richard. What is the least amount of money that must be exchanged so that they each have the same amount of money? A) 4 B) 5 C) 6 D) 7 E) 8

OpenStudy (anonymous):

R = R F = R + 3 G = R +3 + 9 Take 5 from George and give to Richard. Take 2 from George and give to Frank. Now it's: R = R + 5 F = R + 3 + 2 G = R + 3 +9 -7 Which is R = R+5 F = R+5 G = R+5

OpenStudy (anonymous):

Assuming George gives away x dollars, Richard has to receive (12-x) dollars to be even with him, and Frank has to receive 9-x dollars to be even with him. So what is the x for which x=12-x +(9-x)? x= 21 - 2x Solve for x.

OpenStudy (anonymous):

x=7

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

thank you lol